OpenStudy (alones.):

(6 - d)(d^2 - 5 + 3d)

OpenStudy (alones.):

Check my answer

OpenStudy (alones.):

−d^3+3d^2+23^d−30

OpenStudy (alones.):

OOps I mean −d3+3d2+23d−30

OpenStudy (alones.):

@jagr2713 may you check this again?

OpenStudy (alones.):

Yeah i like here more Anita blin udali!!!

OpenStudy (anitas.):

>.< i'ma blin?

OpenStudy (alones.):

mayeb you are -.-

jagr2713 (jagr2713):

Hey, what do you think we do first/

jagr2713 (jagr2713):

Well its correct

OpenStudy (mathmale):

Although I can guess what you're supposed to do here, I'd appreciate your posting instructions with math expressions such as this one. You apparently want to multiply two polynomials together, using the Distributive Property of Multiplication. In the left set of parentheses you have 6-d. Multiply all 3 terms within the right set of parentheses by 6. Next, do the same thing with -3 as your multiplier. Show your work, plese.

OpenStudy (alones.):

We distibute?

OpenStudy (alones.):

Oh okay i get it thanks<3

OpenStudy (kpop4life123):

(6 - d)(d^2 - 5 + 3d) = 6d^2 - 30 + 18d - d^3 + 5d - 3d^2 = -d^3 + 3d^2 + 13d - 30

jagr2713 (jagr2713):

Yes, but you already did it but look at @mathmale work to get better information on it :D

OpenStudy (mathmale):

Your last step would be to "combine like terms."

OpenStudy (alones.):

yeah like this 6d^2−30+18d−d^3+5d−3d^2 and i solve that and get −d^3+3d^2+23d−30

OpenStudy (mathmale):

I appreciate your sharing your work. You've done a nice job of writing out your expression in descending order of the variable d. Strictly speaking, you're not "solving" anything; rather, you're "multiplying two polynomials" or "multiplying a trinomial by a binomial."

OpenStudy (alones.):

Oh okay thanks MM<3